# Burer-Monteiro factorizability of nuclear norm regularized optimization

### v1

@@ Appendix A Pseudo-stationarity
We first consider the case where $G_{1}+\mu X_{1}=G_{2}+\mu X_{2}$. In this case, the function $h$ in A.11 is clearly well defined, and one can verify that the function $h(\cdot)-\frac{\mu}{2}\|\cdot\|_{F}^{2}$ is convex, and $\nabla h$ is Lipschitz continuous with modulus $L$. Moreover, $\nabla h(X_{1})=-G_{1}$ and
$\nabla h(X_{2})=\mu(X_{2}-X_{1})-G_{1}=-G_{2}.$
We start by noticing from 3.2 that $(\bar{U},\bar{V})$ is a stationary point of $F_{r}$ (with $R=I_{m}$, $Q=I_{r}$ and $P=I_{n}$ in 3.2). Consequently, by 3.4, we know that $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$ if and only if for all $U_{11},V_{11}\in\mathbb{R}^{r\times r},~{}U_{21}\in\mathbb{R}^{(m-r)\times r},~{}V_{21}\in\mathbb{R}^{(n-r)\times r}$, [^5] Here, we use the partition in (3.6), which is well defined because $\sigma(\bar{U})=\sigma(\bar{V})$. We also note that $U_{12},V_{12},U_{22},V_{22}$ are void because $\mathrm{rank}(\bar{U})=r$. it holds that
$-2\lambda\mathrm{tr}(U_{11}^{\top}V_{11})-2\mathrm{tr}(D^{\top}U_{21}V_{21}^{\top})+\lambda(\|U_{11}\|_{F}^{2}+\|V_{11}\|_{F}^{2}+\|U_{21}\|_{F}^{2}+\|V_{21}\|_{F}^{2})$ (A.14)
$+\nabla^{2}h(X_{1})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}\geq 0,$
where
$\Sigma=\mathrm{diag}(\sqrt{\sigma_{1}(X_{1})},\dots,\sqrt{\sigma_{r}(X_{1})})\overset{(a)}{=}\mathrm{diag}(\sqrt{v^{1}_{1}},\dots,\sqrt{v^{1}_{r}})\in\mathbb{R}^{r\times r},$ (A.15)
$D=\widetilde{\rm Diag}(d_{r+1}^{1},\dots,d_{m}^{1})\in\mathbb{R}^{(m-r)\times(n-r)},$
and in (a) we have used the fact that $v^{1}$ is a nonnegative vector sorted in descending order. We will verify (A.14).
To this end, we first use the representation of $h$ in A.11 to deduce that
$\begin{split}\nabla^{2}h(X_{1})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}&\geq L\|\Sigma V_{21}^{\top}\|_{F}^{2}+L\|U_{21}\Sigma\|_{F}^{2}\\ &\overset{\rm{(a)}}{\geq}Lv^{1}_{r}(\|U_{21}\|^{2}_{F}+\|V_{21}\|_{F}^{2}),\end{split}$ (A.16)
where in (a) we have used the fact that $\|AB\|_{F}\geq\sigma_{\min}(A)\|B\|_{F}$. Therefore, it holds that
$+\nabla^{2}h(X_{1})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}$
$\overset{\rm{(a)}}{\geq}-2\mathrm{tr}(D^{\top}U_{21}V_{21}^{\top})+\lambda(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+\nabla^{2}h(X_{1})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}$
$\overset{\rm{(b)}}{\geq}-2\|D^{\top}\|_{2}\|U_{21}\|_{F}\|V_{21}^{\top}\|_{F}+\lambda(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+\nabla^{2}h(X_{1})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}$
$\overset{\rm{(c)}}{\geq}-2(\lambda+Lv^{1}_{r})\|U_{21}\|_{F}\|V_{21}\|_{F}+\lambda(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+Lv^{1}_{r}(\|U_{21}\|^{2}_{F}+\|V_{21}\|_{F}^{2})$
$\overset{\rm{(d)}}{\geq}-(\lambda+Lv^{1}_{r})(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+\lambda(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+Lv^{1}_{r}(\|U_{21}\|^{2}_{F}+\|V_{21}\|_{F}^{2})=0,$
where in (a) we have used the Cauchy-Schwartz inequality to show that $\mathrm{tr}(U_{11}^{\top}V_{11})\leq\frac{1}{2}(\|U_{11}\|^{2}_{F}+\|V_{11}\|_{F}^{2})$, in (b) we have used the fact $\mathrm{tr}(ABC)=\mathrm{tr}(CAB)\leq\|C\|_{F}\|AB\|_{F}\leq\|A\|_{2}\|C\|_{F}\|B\|_{F}$, in (c) we have used A.7, A.16 and the definition of $D$ in A.15, and in (d) we have used the fact $\|A\|_{F}\|B\|_{F}\leq\frac{1}{2}(\|A\|_{F}^{2}+\|B\|_{F}^{2})$. This verifies (A.14) and hence $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$.
Next, we consider the case where $G_{1}+\mu X_{1}\neq G_{2}+\mu X_{2}$. By Remark A.4, we know that A.9 is equivalent to
$(L-\mu)\langle(G_{2}+\mu X_{2})-(G_{1}+\mu X_{1}),X_{1}-X_{2}\rangle\geq\|(G_{1}+\mu X_{1})-(G_{2}+\mu X_{2})\|_{F}^{2}.$
Since $G_{1}+\mu X_{1}\neq G_{2}+\mu X_{2}$, we see that $(L-\mu)\langle(G_{2}+\mu X_{2})-(G_{1}+\mu X_{1}),X_{1}-X_{2}\rangle\geq\|(G_{1}+\mu X_{1})-(G_{2}+\mu X_{2})\|_{F}^{2}>0$. In particular, this implies $L>\mu$ and $\langle(G_{2}+\mu X_{2})-(G_{1}+\mu X_{1}),X_{1}-X_{2}\rangle>0$, showing that $h$ in A.10 is well defined. Furthermore, we have
$L-\mu\geq\frac{\|(G_{1}+\mu X_{1})-(G_{2}+\mu X_{2})\|_{F}^{2}}{\langle(G_{2}+\mu X_{2})-(G_{1}+\mu X_{1}),X_{1}-X_{2}\rangle}.$ (A.17)
Now, it is routine to check that $h(\cdot)-\frac{\mu}{2}\|\cdot\|_{F}^{2}$ is convex, $\nabla h(X_{1})=-G_{1}$, and $\nabla h(X_{2})=-G_{2}$. Moreover, the relation in (A.16) and hence the second-order stationarity of $(\bar{U},\bar{V})$ can be verified similarly to that in the case where $G_{1}+\mu X_{1}=G_{2}+\mu X_{2}$. Thus, it remains to show that $\nabla h$ is Lipschitz continuous with modulus $L$.
To this end, notice that for all $X,Y\in\mathbb{R}^{m\times n}$ and the function $h$ defined in A.10, it holds that
$\nabla^{2}h(X)[Y,Y]=L\sum_{i=1}^{m}\sum_{j\neq i}^{n}Y_{ij}^{2}+\mu\sum_{i=1}^{n}Y_{ii}^{2}+\frac{(\langle Y,-G_{2}-\mu X_{2}+G_{1}+\mu X_{1}\rangle)^{2}}{\langle X_{2}-X_{1},-G_{2}-\mu X_{2}+G_{1}+\mu X_{1}\rangle}$
=$\overset{\rm{(a)}}{=}L\sum_{i=1}^{m}\sum_{j\neq i}^{n}Y_{ij}^{2}+\mu\sum_{i=1}^{n}Y_{ii}^{2}+\frac{(\langle\widetilde{\rm Diag}(Y_{11},\dots,Y_{mm}),-G_{2}-\mu X_{2}+G_{1}+\mu X_{1}\rangle)^{2}}{\langle X_{2}-X_{1},-G_{2}-\mu X_{2}+G_{1}+\mu X_{1}\rangle}$
$\overset{\rm(b)}{\leq}L\sum_{i=1}^{m}\sum_{j\neq i}^{n}Y_{ij}^{2}+\mu\sum_{i=1}^{n}Y_{ii}^{2}+\frac{\|\widetilde{\rm Diag}(Y_{11},\dots,Y_{mm})\|^{2}_{F}\,\|-G_{2}-\mu X_{2}+G_{1}+\mu X_{1}\|_{F}^{2}}{\langle X_{2}-X_{1},-G_{2}-\mu X_{2}+G_{1}+\mu X_{1}\rangle}\overset{\rm(c)}{\leq}L\|Y\|_{F}^{2},$
where in (a) we have used the fact that $X_{1},X_{2},G_{1},G_{2}$ are diagonal (see A.8), in (b) we have used the Cauchy-Schwartz inequality, and in (c) we have used (A.17). This proves that $\nabla h$ is Lipschitz continuous with modulus $L$, and completes the proof. ∎
Appendix B Proof of 5.2
This subsection contains the essential auxiliary results for the proof of 5.2. Note that 5.2 is about deriving conditions for $\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$ to consist solely of $h$ whose corresponding $f$ in 1.4 is $r$-factorizable. Our first task is to reduce this problem to a simpler one.
Proposition B.1.
Let $\infty>L\geq\mu>0$, $r\in[m]$ and $r^{*}\in[m]\cup\{0\}$. The following statements are equivalent.
  1. (i)
    There exists an $h\in\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$ (see 5.1) such that $f$ in 1.4 is not $r$-factorizable.
  2. (ii)
@@
    There exist $x,g,y,v\in\mathbb{R}^{m}$ with $\|g\|_{\infty}>\lambda$ and $\tau\in\mathfrak{P}_{m}$ such that
    $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (B.1a)
@@
    $\forall i\in[r],~{}x_{i}>0,~{}g_{i}=\lambda,\quad\forall i\in[r^{*}],~{}y_{i}>0,~{}v_{i}=\lambda,$ (B.1b)
    $\forall i\in[m]\setminus[r],~{}x_{i}=0,~{}g_{i}\in[0,\lambda+L\min_{j\in[r]}x_{j}],\quad\forall i\in[m]\setminus[r^{*}],~{}y_{i}=0,~{}v_{i}\in[0,\lambda].$ (B.1c)
@@
Assume (i). By (i), we can select $h\in\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$ and $X_{2}\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(X_{2})=r^{*}$, $-\nabla h(X_{2})\in\lambda\partial\|X_{2}\|_{*}$, and $f$ is not $r$-factorizable. The latter means we can find $(\bar{U},\bar{V})$ being a second-order stationary point of $F_{r}$ in 1.5 and $X_{1}=\bar{U}\bar{V}^{\top}$ is not a stationary point of $f$.
Applying 3.2 (see also Remark 3.3) to $(\bar{U},\bar{V})$ and using 2.1 and the condition that $-\nabla h(X_{2})\in\lambda\partial\|X_{2}\|_{*}$, we can write
$X_{1}=R_{1}\begin{bmatrix}\Sigma_{1}&0\end{bmatrix}P_{1}^{\top},~{}-\nabla h(X_{1})=R_{1}\begin{bmatrix}D_{1}&0\end{bmatrix}P_{1}^{\top},$ (B.2)
@@
for some $R_{i}\in\mathcal{O}^{m}$ and $P_{i}\in\mathcal{O}^{n}$ and $m\times m$ diagonal matrices $\Sigma_{i}$ and $D_{i}$, $i=1,2$, where $\mathrm{diag}(\Sigma_{i})\in\mathbb{R}^{m}_{+}$ consisting of all the singular values of $X_{i}$ in descending order, $d^{i}:=\mathrm{diag}(D_{i})\in\mathbb{R}^{m}_{+}$ with $d^{i}_{1}=\dots=d^{i}_{\mathrm{rank}(X_{i})}=\lambda$, for $i=1,2$. Clearly, we have $\mathrm{rank}(X_{1})=r$, otherwise by 4.4 we can conclude that $X_{1}$ is a stationary point of $f$, leading to a contradiction. Applying Lemma A.2, there exists $\bar{\tau}\in\mathfrak{P}_{m}$ such that
$\sum_{i=1}^{m}(L\sigma_{i}(X_{1})+d^{1}_{i})(\mu\sigma_{\bar{\tau}(i)}(X_{2})+d^{2}_{\bar{\tau}(i)})+\sum_{i=1}^{m}(\mu\sigma_{i}(X_{1})+d^{1}_{i})(L\sigma_{\bar{\tau}(i)}(X_{2})+d^{2}_{\bar{\tau}(i)})$ (B.3)

### v2

@@ 3 First- and second-order stationary points of $F_{r}$
A pair $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ is a second-order stationary point of $F_{r}$ if and only if it satisfies both the first- and second-order optimality conditions. By Proposition 3.2, the first-order optimality condition is equivalent to (i).
Using 3.13, the second-order condition in 3.12 can be rewritten as that for all $(U,V)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$,
where (a) follows from the decomposition for $\nabla h(\bar{X})$ in (i) and the definition of $D$ and $\Sigma$, and (b) from the definition of the blocks in 3.14. This shows that (i) and (ii) together form an equivalent characterization of the second-order stationary points.
We next prove the second claim. Assume that $s=\mathrm{rank}(\bar{U})<r$. Note that this implies that $U_{22}$ and $V_{22}$ are not null. Hence, we can take $U$ and $V$ in 3.15 to be the matrices with the blocks $U_{22}=[e_{1}~0]$ and $V_{22}=[e_{1}~0]$ and $U_{11}$, $U_{12}$, $U_{21}$, $V_{11}$, $V_{12}$, $V_{21}$ all being zero matrices to deduce that
$2\sigma_{1}(D)+2\lambda\geq 0.$
The desired conclusion now follows immediately from the above display and the decomposition of $\nabla h(\bar{U}\bar{V}^{\top})$ in (i). ∎
Using the second-order condition in Proposition 3.6, we now derive the information on the singular values of $\nabla h(\bar{X})$, when $\bar{X}=\bar{U}\bar{V}^{\top}$ with $(\bar{U},\bar{V})$ being a second-order stationary point of $F_{r}$ in 1.2. The technique used here is not new, and similar arguments can be found in [11] and [3].
Theorem 3.7 (Singular values for 2nd stationary points).
Let $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ be a second-order stationary point of $F_{r}$ in 1.2, $s=\mathrm{rank}(\bar{U})$ and $d\in\mathbb{R}^{m}$ be given in Proposition 3.6. Then, [^7] We define $d_{m+1}=0$. $d_{s+1}\leq\lambda+\tilde{L}\sigma_{r}(\bar{U}\bar{V}^{\top})$, where
$\tilde{L}:=\sup_{\begin{subarray}{c}Y\in\mathbb{R}^{m\times n},\,\|Y\|_{F}=1,\\ \mathrm{rank}(Y)=2\end{subarray}}\nabla^{2}h(\bar{U}\bar{V}^{\top})[Y,Y].$
Proof.
$d_{s+1}\leq\|d\|_{\infty}=\|\nabla h(\bar{U}\bar{V}^{\top})\|_{2}\leq\lambda=\lambda+\tilde{L}\sigma_{r}(\bar{U}\bar{V}^{\top}),$
where the first equality follows from Remark 3.3(i) and the second inequality follows from Proposition 3.6.
We next consider the case where $\mathrm{rank}(\bar{U})=r$. If $r=m$, then $d_{m+1}=0$ and the desired conclusions then hold trivially. Thus, from now on, we assume $r<m$.
By Proposition 3.2, $\mathrm{rank}(\bar{V})=\mathrm{rank}(\bar{U})=r$. Therefore, in this case, the blocks $U_{12},U_{22},V_{12},V_{22}$ in (3.14) are null. Since $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$ in 1.2, by Proposition 3.6, it satisfies the following inequality for any matrices $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{n\times r}$:
Taking $U$ and $V$ to be the matrices with $U_{11}$ and $V_{11}$ being zero and $U_{21}=[e_{r}\ 0]^{\top}=e_{1}e_{r}^{\top}\in\mathbb{R}^{(m-r)\times r}$ and $V_{21}=[e_{r}\ 0]^{\top}=e_{1}e_{r}^{\top}\in\mathbb{R}^{(n-r)\times r}$, we have that $\Sigma V_{21}^{\top}=\sigma_{r}(\bar{U})e_{r}e_{1}^{\top}$ and $U_{21}\Sigma=\sigma_{r}(\bar{U})e_{1}e_{r}^{\top}$ and that $\mathrm{tr}(D^{\top}U_{21}V_{21}^{\top})=e_{1}^{\top}De_{1}=d_{r+1}$. Substituting these into (3.18) yields
[…]
where the second inequality follows from the fact that $\left\|\sigma_{r}(\bar{U})R\begin{bmatrix}0_{r\times r}&e_{r}e_{1}^{\top}\\ e_{1}e_{r}^{\top}&0_{(m-r)\times(n-r)}\end{bmatrix}P^{\top}\right\|_{F}=\sqrt{2}\sigma_{r}(\bar{U})$ and the definition of $\tilde{L}$, and the equality follows from Remark 3.3(ii). Hence, $d_{r+1}\leq\lambda+\tilde{L}\,\sigma_{r}(\bar{U}\bar{V}^{\top})$. ∎
In view of 3.7, the definition of $d$ in Proposition 3.6 and Proposition 2.1, for a second-order stationary point $(\bar{U},\bar{V})$ of $F_{r}$ in 1.2, by setting $\bar{X}=\bar{U}\bar{V}^{\top}$, we see that if $\sigma_{r}(\bar{X})=0$, then $\bar{X}$ is a stationary point of $f$, as proved in [3]; see, also [32, Section 3], and [15, 4, 5] for similar results. This is formally presented as the next corollary.
Corollary 3.8.
Intuitively, based on the bound on $d_{s+1}$ in 3.7, we can say that the smaller $\sigma_{r}(\bar{X})$ is, the closer $\bar{X}$ is to being a stationary point of $f$ in 1.1.
4 Characterization of $r$-factorizability for strongly convex functions
Suppose we have an $h\in\mathfrak{S}(L,\mu,r^{*})$ in hand such that the corresponding $f$ in 1.1 is not $r$-factorizable. By definition, we know there exists $X^{*}\in\mathbb{R}^{m\times n}$ such that $X^{*}$ is a stationary point of $f$. Let $(\bar{U},\bar{V})$ be the second-order stationary point of $F_{r}$ in 1.1 such that $\bar{X}=\bar{U}\bar{V}^{\top}$ is not a global minimizer of $f$ (or equivalently, not a stationary point of $f$ since $f$ is strongly convex thanks to $h\in\mathfrak{S}(L,\mu,r^{*})$). Then according to Proposition 3.2, we know $\bar{X}$ is a pseudo stationary point of $f$. In this case, we have the bound 3.7 on the singular values of $X^{*}$, $\nabla h(X^{*})$, $\bar{X}$ and $\nabla h(\bar{X})$. In addition, for a pseudo stationary point $X$ of $f$, it is a stationary point of $f$ if and only if $\|\nabla h(X)\|_{2}\leq\lambda$ in view of Definition 3.4 and Proposition 2.1. Finally, the singular value of $\nabla h(\bar{X})$ is further constrained by 3.7. Putting all these conditions together, we have the following result.
Proposition 4.1.
@@ 4 Characterization of $r$-factorizability for strongly convex functions
Let $L\geq\mu>0$, $r\in[m]$ and $r^{*}\in[m]\cup\{0\}$. Assume that there exists an $h\in\mathfrak{S}(L,\mu,r^{*})$ (see Definition 1.2) such that $f$ in 1.1 is not $r$-factorizable. Then, there exist $x,g,y,v\in\mathbb{R}^{m}$ with $\|g\|_{\infty}>\lambda$ and $\tau\in\mathfrak{P}_{m}$ such that
$\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (4.1a)
@@
$\forall i\in[r],~x_{i}>0,~g_{i}=\lambda,\quad\forall i\in[r^{*}],~y_{i}>0,~v_{i}=\lambda,$ (4.1b)
$\forall i\in[m]\setminus[r],~x_{i}=0,~g_{i}\in[0,\lambda+L\min_{j\in[r]}x_{j}],\quad\forall i\in[m]\setminus[r^{*}],~y_{i}=0,~v_{i}\in[0,\lambda].$ (4.1c)
@@
By assumption, we can select $h\in\mathfrak{S}(L,\mu,r^{*})$ and $X_{2}\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(X_{2})=r^{*}$, $\nabla h(X_{2})\in\lambda\partial\|X_{2}\|_{*}$, and $f$ is not $r$-factorizable. The latter means we can find $(\bar{U},\bar{V})$ being a second-order stationary point of $F_{r}$ in 1.2 and $X_{1}=\bar{U}\bar{V}^{\top}$ is not a stationary point of $f$.
Applying Proposition 3.2 (see also Remark 3.3) to $(\bar{U},\bar{V})$ and using Proposition 2.1 and the condition that $\nabla h(X_{2})\in\lambda\partial\|X_{2}\|_{*}$, we can write
$X_{1}=R_{1}\begin{bmatrix}\Sigma_{1}&0\end{bmatrix}P_{1}^{\top},~-\nabla h(X_{1})=R_{1}\begin{bmatrix}D_{1}&0\end{bmatrix}P_{1}^{\top},$ (4.2)
@@
for some $R_{i}\in\mathcal{O}^{m}$ and $P_{i}\in\mathcal{O}^{n}$ and $m\times m$ diagonal matrices $\Sigma_{i}$ and $D_{i}$, $i=1,2$, where $\mathrm{diag}(\Sigma_{i})\in\mathbb{R}^{m}_{+}$ consisting of all the singular values of $X_{i}$ in descending order, $d^{i}:=\mathrm{diag}(D_{i})\in\mathbb{R}^{m}_{+}$ with $d^{i}_{1}=\dots=d^{i}_{\mathrm{rank}(X_{i})}=\lambda$, for $i=1,2$. Clearly, we have $\mathrm{rank}(X_{1})=r$, otherwise by Corollary 3.8 we can conclude that $X_{1}$ is a stationary point of $f$, leading to a contradiction. Applying Lemma 3.5, there exists $\bar{\tau}\in\mathfrak{P}_{m}$ such that
[…]